OpenStudy (zeesbrat3):

A girl is floating in a freshwater lake with her head just above the water. If she weighs 610 N, what is the volume in cubic meters of the submerged part of her body?

OpenStudy (zeesbrat3):

Any ideas?

OpenStudy (zeesbrat3):

it is a short response question.

OpenStudy (michele_laino):

If I call with V the requested volume, then I can write: \[\large V \times {\rho _{{H_2}O}} \times g = P\] where: \[\large g = 9.81\;m/{\sec ^2}\] \[\large {\rho _{{H_2}O}}\] is the density of water, and P=610 N

OpenStudy (zeesbrat3):

and the density of water would be 999.97 kg/m^3, if that has any significance here...

OpenStudy (michele_laino):

approximately 1,000 Kg/m^3

OpenStudy (zeesbrat3):

yes. so v x 999.97 x 9.81 = 610?

OpenStudy (michele_laino):

that's right!

OpenStudy (zeesbrat3):

16.1?

OpenStudy (michele_laino):

I got 0.0621 m^3

OpenStudy (zeesbrat3):

whoops, divided the wrong side!

OpenStudy (zeesbrat3):

can you help me with one more?

OpenStudy (michele_laino):

yes!

OpenStudy (zeesbrat3):

Huckleberry Finn's wooden raft measuring 4.9 m by 3.7 m by 0.36 m had a mass of 977 kg. If the raft is ti stay afloat in fresh water, what is the maximum amount of extra mass in kilogram that can be safely added?

OpenStudy (zeesbrat3):

@Michele_Laino

OpenStudy (michele_laino):

hint: here we can write the subsequent equation (Archimede's principle): \[\large \left( {{m_{add}} + {m_{raft}}} \right) \times g = V \times {\rho _{{H_2}O}} \times g\] where: \[{{m_{add}}}\] is the added mass

OpenStudy (michele_laino):

and: \[{{m_{raft}}}\] is the mass of the raft

OpenStudy (zeesbrat3):

\[(m _{add} + 977) \times 9.81 = 6.5268 \times 999.97 \times 9.81\]

OpenStudy (michele_laino):

and: \[\large V = 4.9 \times 3.7 \times 0.36 = ...?\]

OpenStudy (michele_laino):

that's right!

OpenStudy (zeesbrat3):

\[m_{add} = 5549.6\]

OpenStudy (michele_laino):

that's right!

OpenStudy (zeesbrat3):

Thank you!

OpenStudy (michele_laino):

Thank you!